OpenStudy (anonymous):

line has an x-intercept of 6 and a y-intercept of 3. write an equation in standard form.

OpenStudy (anonymous):

NCA 9th Grade Algebra 1A U5 L8

OpenStudy (aripotta):

ok so x-intercept is 6 which gives us the point (6,0) and the y-intercept is 3 which gives us the point (0,3) from those two points, we can find the slope of the line using the slope formula:\[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\] once you find the slope, you can use one of the points and put the info into point-slope form:\[y - y _{1}=m(x - x _{1})\] then you can convert from point-slope form to standard form, which is ax + by = c